OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

let sin x = x

OpenStudy (ghazi):

again consider sin^2x= y and solve for y

OpenStudy (anonymous):

just treat it as a quadratic equation...

OpenStudy (anonymous):

when sin x = x then u get 7x^2-14x +7 = 0

OpenStudy (anonymous):

I'm pretty sure there is 4 solutions

OpenStudy (anonymous):

2sinxcosx=-2^1/2cosx cosx=0 x=90 & 270degrees sinx=-2^2/2 x=-45,+45degrees

OpenStudy (ghazi):

yes there are four solution for this

OpenStudy (anonymous):

then from here i guess u should know how to factorize and solve for x rite? if u dont know then.. here is the solution.. 7(x^2-2x+1) = 0

OpenStudy (anonymous):

(x-1)(x-1) = 0 x = 1... so sin x = 1 when is sin x = 1? when x = pi/2

OpenStudy (anonymous):

and the nxt solution is pi/2 + 2pi

OpenStudy (anonymous):

3pi/2

OpenStudy (anonymous):

yeap..

OpenStudy (anonymous):

5pi/2?

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

eh wait.. 3pi on 2 is not

OpenStudy (anonymous):

3pi/2 is 270 degrees it will give u sin x = -1 not 1

OpenStudy (anonymous):

wait so what are the four factors?

OpenStudy (anonymous):

sorry 2kpi +pi/2

OpenStudy (ghazi):

(2n*pi)+pi/2

OpenStudy (anonymous):

@babydoll332 four factors?

OpenStudy (anonymous):

just kidding so all i can do is say (2n*pi)+pi/2

OpenStudy (anonymous):

yes,but you must put,that n lies in integers

OpenStudy (anonymous):

which means?

OpenStudy (ghazi):

n= 1,2,3...

OpenStudy (anonymous):

ohh ok

OpenStudy (ghazi):

also 0

OpenStudy (anonymous):

n=...,-1,0,1.....

OpenStudy (anonymous):

negative too

OpenStudy (anonymous):

thanks for helping me everyone :)